OpenStudy (anonymous):

Suppose that you find in a reference book that the volume of all the oceans is 1.4×109km3. To find the mass, you can use the density of water, also found in this reference book, but first you must convert the volume to cubic meters. What is this volume in cubic meters? Express your answer in cubic meters.

OpenStudy (ybarrap):

\( 1~km^3 = (10^3~ m)^3=10^9m^3\), which is 1 billion cubic meters \(v=1.4×109~km^3\times \dfrac {10^9m^3}{1~km^3}=1.4\times109\times10^9m^3 \) \(\implies v=1.526\times10^{11}~m^3\)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

especially showing the steps

OpenStudy (anonymous):

what is the exponet you have listed 10^?m^2 Im not understanding how you got 10^11